++++ INAF LBC+LUCIFER queue observing, 2011 March 28 - 29

Observer: R. Speziali, L. Pentericci, K. Boutsia
Telescope Operator: G. Becchetti
Support Astronomer: O. Kuhn
Instrument: LBC - LUCIFER

Summary

Clear night with seeing <1.0'' - could be photometric
The Buttiglione program has been repeated and now it is within specs - completed
We had problems with Lucifer - first the compressor alarm went off and it was giving the INSTRUMENT CRITICAL error on the intrument panel
We were not able to center the HDF_K1 mask of Maiolino because the MOS ACQ routine was not giving the correct offsets

Data Summary

Project Objects Filter ExpTime Notes
Buttiglione HATLAS B-bessel, z-sloan 1.5h seeing ~1.0'' - within specs this time
Grazian EGS U-spec, r-sloan 20min seeing >1.2'' - OB Aborted
Gallerani SDSSJ1048+4637_ref 210zJHK+Z 2hr seeing measure on the image <1.0'' - Tell +ARC+FLAT
PHOT_STD SA098 UsBg_rz    
PHOT_STD SA101 UsBg_rz    

Details

Begin night with LBC

01:27 -Connect to LBT - 10Bias_Bino
  • There is a gradient in the biases - red channel - left to right
  • Biases taken with the dome open!!

01:56 - Skyflat
  • Uspec - z : 2sec - 0.5sec PA=0
  • Bbess - z : 1.1sec - 1sec PA=0
  • Bbess - z : 3.3sec - 3.0 sec PA=180 e le ultime 2 esposizioni 6.6 sec e 6.0
  • g-sloan - r-sloan : 5.0sec - 6.0 sec PA=180
  • g-sloan - r-sloan : 15.0sec - 18.0 sec PA=0

02:25 - Slew to pointing star WT10_197 * dofpia before correcting pointing - almost converged - calculated seeing Blue 1.9/Red 0.97 - 1 deg tem difference for the mirrors * Pointing correction OK + mirror adjust OK with the new IRAF script lbcrangebal

02:41 - slew to PHOT_STD - SA098 dofpia,
  • we manually gave positive Z22 at the left side because pupils looks fuzzy.
  • it took us some time to converge but the results seems much better than yesterday.
  • when started the exposure airmass of the standard is ~1.2
  • dimm gives ~1.0'' seeing

03:13 - slew to Buttiglione_HATLAS_1 -dofpia

03:21 - observing target Buttiglione_HATLAS_1
  • image quality much better
  • seeing measured on the images is <1.0'' in both sides

04:05 - dofpia - 2 iterations to converge

04:09 - observing target Buttiglione_HATLAS_2
  • seeing on the dimm 0.85''

04:52 - dofpia

04:56 - observing target Buttiglione_HATLAS_3
  • seeing 0.9'' - 1.1''

05:39 - Slew to PHOT_STD - SA101 dofpia
  • filters observed Uspec, Bbess, g-sloan_z-sloan, r-sloan

05:49 - Slew to Grazian EGS - dofpia
  • it took 4 iterations to converge
  • seeing seems to get worse - we are measuring ~1.2''

05:59 - grazian_EGS_UR_1
  • OB aborted after 20 minutes because seeing is getting worse and images are out of spec
  • dimm 1.2-1.3'' and seeing measured on the blue images is >1.3''

06:20 - SWITCH TO LUCIFER

06:55 - Slew to Gallerani SDSSJ1048+4637_ref for acq 0002-0005
  • centering on the target and not the reference star

07:17 - SG9_spec_science_Z_3.txt 0006-0017
  • seeing measured on the images <1.0''
  • seeing on the dimm is variable 0.7''-1.2''

08:25 - SG9_spec_science_Z_4.txt 0018-0029
  • first we made an Y-9.5'' offset to position object at the initial position
  • seeing measured on the image 0.9''
  • there is an alarm - INSTRUMENT CRITICAL! - on the instrument control GUI
  • we go on with the OB

09:31 - Tell acq for Gallerani 0030-0033

09:41 - Telluric science script for Gallerani 0034-0036
  • dimm seeing 1.2''

09:45 - Calib Unit IN
  • ARC 0037-0040
  • FLAT 0041-0052

09:55 - Finished with Calibration
  • there is a compressor alarm
  • bring telescope to zenith in order to manually reset the compressor alarm
  • after 10 minutes they decide not to intervene - we can go on with the acquisition of the next field

10:04 - MOS acquisition for Maiolino HDF_K1 0053-0055
  • the mask pin-holes look elongated and we INITIALIZE 0056
  • the problem persists and we initialize again 0057
  • the second time it was a bit better so we switch from N3.75 to N1.8 and then initialize for the 3rd time
  • this times the pin-holes look sharp - we start all over again 0059-0061
  • the MOS ACQ button gives almost 0.0 offsets
  • we don't see the objects and we increase the exposure time to 25x6 0062-0067
  • after the last MOS ACQ the upper object seems to be in the slit, but the bottom object seems to be offset - is there an error in calculating the rotation (??)

12:07 - twilight is passed and we still don't understand why the MOS ACQ does not calculate the correct offsets
  • Olga is repeating the acquisition trying to figure out what went wrong 0068-0071

12:35 - EndofNight Script - DoDarks

-- KonstantinaBoutsia - 28 Mar 2011
Topic revision: r6 - 31 Mar 2011, KonstantinaBoutsia
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback